OpenStudy (anonymous):

Brookner has 3 pipes. One can fill a pool in 3 hours. Another can drain the same pool in 4 hours. The last one can fill the same pool in 9 hours. If all three pipes are left open, how long will it take, in hours, to fi ll an empty pool?

OpenStudy (whpalmer4):

First pipe has a rate of 1 pool / 3 hours Second pipe has a rate of -1 pool / 4 hours Third pipe has a rate of 1 pool / 9 hours Add up the rates, find the time to fill 1 pool.

OpenStudy (anonymous):

\[1/3 -1/4 +1/9\]=7/36 7x/36 = 1 x = 36/7 is that correct?

OpenStudy (anonymous):

@whpalmer4

OpenStudy (whpalmer4):

That is correct.